At a certain college, it is estimated that fewer than 75% of the students have cars on campus, and with this number, the parking lot is fully occupied. Would you suggest to increase the parking lot space if, in a random sample of 60 college students, 55 are found to have cars? Use 0.1 level of significance.


1
Expert's answer
2022-04-16T04:13:44-0400

With given level of significance real proportion of students having cars is bigger than

"\frac{55}{60}-0.1\frac{55}{60}=0.9\frac{55}{60}=\\
=0.825=82.5\%\gt75\%,"

so I would suggest to increase the parking lot space.